Please help me solve this equation tutor 3x+y+13 and 3x+2y=17 using substitution
Answer by tutorcecilia(2152)

3x+y=13 and 3x+2y=17
3x+y=13 [Pick one equation and solve for y (or x)]
3x-3x+y=13-3x
y=13-3x [Plug this back into the second equation for the "y"]
.
3x+2(13-3x)=17 [Simplify]
3x+26-6x = 17
-3x+26=17 [Solve for x]
-3x+26-26=17-26
-3x=-9
-3x/-3=-9/-3
x=3
.
3x+y=13 and 3x+2y=17
[Plug (x=3) back into the original equations to solve for y]
.
3x+y=13
3(3)+y=13
9+y=13
y=13-9
y=4
.
and
.
3x+2y=17
3(3)+2y=17
9+2y=17
2y=8
y=4 [Also, checks out]
.
So (x=3) and (y=4)
3(3)+(4)=13
9+4=13
13=13 [checks out}
and
3(3)+2(4)=17
9+8=17
17=17 [Also checks out]